OpenStudy (anonymous):

The mean weight of four members of the wrestling team is 155 pounds. If three of them have weights of 125, 183, and 169 pounds, then what does the fourth wrestler weigh? 127 pounds 143 pounds 155 pounds 158 pounds

OpenStudy (auctoratrox):

(125+183+169+x)/4 = 155

OpenStudy (anonymous):

\[(125+183+169+x)/4=155\]\[(477+x)=620\]\[143\]